.     Record the following transactions in the financial statements effects template below.

  1. Founder contributes $44,000 in cash in exchange for common stock.
  2. Obtain $26,000 short-term bank loan.
  3. Purchase equipment costing $24,000 for cash.
  4. Purchase inventory costing $14,000 on account.

Balance Sheet

Income Statement

Transaction

Cash Asset

+

Noncash Assets

=

Liabil-

ities

+

Contrib. Capital

+

Earned

Capital

Rev-enues

Expen-ses

=

Net

Income

a)

=

=

b)

=

=

c)

=

=

d)

=

=
Answer #1

Following is the financial statements effect template as updated by transactions given :-

Transaction balance sheet Income statement

Cash

Asset

+

Non Cash

Asset

= liabilities + contributed capital + earned capital revenues - Expenses = net income e
a) . 44000 + 0 = 0 + 44000 + 0 0 - 0 = 0
b) . 26000 + 0 = 26000 + 0 + 0 0 - 0 = 0
c) . (24000) + 24000 = 0 + 0 + 0 0 - 0 = 0
d) . 0 + 14000 = 14000 + 0 + 0 0 - 0 = 0